OpenStudy (anonymous):

What is the radius of a circle in which a 30° arc is 2 inches long?

OpenStudy (anonymous):

Hi Jenny

OpenStudy (anonymous):

That I dont know...

OpenStudy (anonymous):

Use proportions to solve \[\frac{ 30 }{ 360 }=\frac{ 2 }{ 2\pi r }\] Solve for r.